OpenStudy (anonymous):

Solve using the elimination method. Show the work. Does the system have no solution or an infinite number of solutions? x + y = 10, x – y = 2

OpenStudy (chaise):

x+y=10 x-y=2 Take the two equations away from one another, remembering a positive minus a negative is a positive. x+y=10 x-y=2 2y=8 y=4 Substitute this back into an original equation. x+y=10 x+4=10 x=6 x-y=2 x-4=2 x=6

OpenStudy (chaise):

There is only one set of two possible solutions, one for x and one for y.

OpenStudy (anonymous):

I think my answer is x=6 and y=4. Thanks!

OpenStudy (anonymous):

I'm attempting to work on the problem(s) as I post them. Thanks! This reinforces concepts learned years ago:)